OpenStudy (anonymous):

If P pencils cost D dollars, how many pencils can be bought for C cents? I said d/p=100c/x so x(d/p)=100c x=100pc/d but the answer is pc/100d What did I do wrong?

OpenStudy (allank):

Let me write the methodology to the answer then compare it to yours. P pencils cost D dollars = D*100 cents Thus one pencil costs 100d/p. The number of pencils that can be bought by c cents will be c/(100d/p) = pc/100d

OpenStudy (anonymous):

perfect why do you multiply d by 100 100cent=1 dollar 100c=1d

OpenStudy (allank):

I do that because c and d are unknowns. d may be 5 dollars. To convert 5 dollars into cents, we multiply by 100, right?

OpenStudy (anonymous):

I understand what you did but when I look at my work I can't see the mistake. I know I did something wrong but what is it?

OpenStudy (allank):

Your first statement is what made your working incorrect. d/p=100c/x Let's interpret it. d/p gives is the price of a pencil in dollars, right?

OpenStudy (anonymous):

yup

OpenStudy (allank):

Good. Now, 100c is 100*(a certain number of cents) right?

OpenStudy (anonymous):

yes to change cents to dollars

OpenStudy (allank):

Remember, to get the number of dollars, we should divide by 100. not multiply. :)

OpenStudy (anonymous):

ohhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhh I kept thinking 100cents=1 dollar so you multiply

OpenStudy (anonymous):

thanks I get stupid sometimes :D

OpenStudy (allank):

Glad to be of help :)